OpenStudy (anonymous):

what is the area of a regular 15-gon with a perimeter of 90 m ?

OpenStudy (anonymous):

Assuming a regular polygon, then s =90/15=6 A 15- gon has exterior angles = 360/15=24, so interior angles of 156 Form a triangle with a radius, apothem and a half- side Then the 156/2= 78 degrees, with the apothem being opposite, and the halfside being adjacent= 3 so tan 78= a/3 and a= 3tan 78=14.114 Then area = (1/2)Pa= (1/2)(90)(a)=635.13 m^2

OpenStudy (anonymous):

ok this makes sense, I just need a proper calculator to perform the function. but thank you, I will remember that, I have the similar example in my book, just a bit tricky.

OpenStudy (anonymous):

Ok...I hope that helps!

OpenStudy (anonymous):

it definitely does, thank you!